Apply the method of variations of parameters to solve the following differential equations:
a) x^2y''+xy'-y=x^2e^x
b)y"+a^2y=cosecax
1
Expert's answer
2020-09-14T15:35:01-0400

a). At first, we solve the following differential equation:

"x^2y''+xy'-y=0"

We look for solution in the following form:

"y=x^{\alpha}."

After substitution we receive:

"\alpha(\alpha-1)x^{\alpha}+\alpha x^{\alpha}-x^{\alpha}=0"

From this we receive:

"\alpha^2-1=0."

Thus, we have the solution "y=c_1x+c_2\frac{1}{x}" , "c_1,c_2\in\mathbb{C}" .

Now, we set "y=c_1(x)x+c_2(x)\frac{1}{x}" and add additional restriction:

"c_1'x+c_2'\frac1x=0". The latter implies "c_1''x+c_1'+c_2''\frac1{x}-c_2'\frac1{x^2}=0" . We substitute "y" into "x^2y''+xy'-y=x^2e^x"

and get:

"x^2(c_1''x+2c_1'+c_2''\frac1x-2c_2'\frac1{x^2}+2c_2\frac1{x^3})+x(c_1'x+c_1+c_2'\frac1{x}-c_2\frac{1}{x^2})-(c_1x+c_2\frac{1}{x})=x^2e^x"

"x^2(c_1'-c_2'\frac1{x^2})=x^2e^x"

Thus, we receive a system of equations:

"x^2(c_1'-c_2'\frac1{x^2})=x^2e^x"

"c_1'x+c_2'\frac1x=0"

From the latter we obtain:

"2c_1'x=xe^x"

Thus, "c_1(x)=\frac12e^x+k_1" , "c_2(x)=-\frac12(2-2x+x^2)e^x+k_2" , "k_1,k_2\in\mathbb{C}"

The resulting solution is "y=(\frac12e^x+k_1)x+(-\frac12(2-2x+x^2)e^x+k_2)\frac{1}{x}" .

b). At first, we solve the equation: "y''+a^2y=0" . The solution is "y=c_1e^{i a x}+c_2e^{-i a x}"

We suppose that "y=c_1(x)e^{i a x}+c_2(x)e^{-i a x}" and "c_1'e^{i a x}+c_2'e^{-i a x}=0" . From the latter we get: "c_1''e^{i a x}+iac_1'e^{iax}+c_2''e^{-i a x}-iac_2'e^{-iax}=0" We substitute "y" into "y''+a^2y=\frac{1}{sin \,ax}" and receive "c_1''e^{i a x}+2i\,a\,c_1'e^{i a x}-c_1a^2e^{i a x}+c_2''e^{-i a x}-2iac_2'e^{-i a x}-c_2a^2e^{-i a x}+a^2(c_1(x)e^{i a x}+c_2(x)e^{-i a x})=\frac{1}{sin \,ax}"

"i\,a\,c_1'e^{i a x}-iac_2'e^{-i a x}=\frac{1}{sin \,ax}"

"c_1'e^{i a x}+c_2'e^{-i a x}=0"

"2i\,a\,c_1'e^{i a x}=\frac{1}{sin \,ax}"

"c_1'=-i\frac{e^{-i a x}}{2a\,sin \,ax}"

"c_1=-\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)-ix+c)"

"c_2=\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)+ix+c)"

"y=-\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)-ix+c)e^{i a x}+\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)+ix+c)e^{-i a x}"

Answer: a) "y=(\frac12e^x+k_1)x+(-\frac12(2-2x+x^2)e^x+k_2)\frac{1}{x}"

b) "y=-\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)-ix+c)e^{i a x}+\frac{i}{2a}(\frac{1}{a}ln(sin\,ax)+ix+c)e^{-i a x}"